# The Klein-Gordon equation on asymptotically Minkowski spacetimes: causal propagators

### v1

@@ 6. Propagation estimates over $C$
$\lvert\operatorname{Im}\langle Q_{r}P_{V}u,Q_{r}u\rangle\rvert\geq-\frac{1}{4\mu}\lVert{}^{\mathrm{sc}}\Lambda_{-1/2,1/2}Q_{r}P_{V}u\rVert^{2}-\mu\lVert{}^{\mathrm{sc}}\Lambda_{1/2,-1/2}Q_{r}u\rVert^{2}\,.$
Since ${}^{\mathrm{sc}}\Lambda_{1/2,-1/2}Q_{r}$ is a 0th order multiple of $B_{r}$ and ${}^{\mathrm{sc}}\Lambda_{-1/2,1/2}$ is a 0th order multiple of $Q_{1,r}$, we can absorb $\mu\lVert{}^{\mathrm{sc}}\Lambda_{1/2,-1/2}Q_{r}u\rVert^{2}$ into $\frac{1}{2}\lVert B_{r}u\rVert^{2}$ if we choose $\mu\in(0,1)$ small enough modulo lower order terms. Then, we conclude that
$\lVert B_{r}u\rVert^{2}\lesssim\lVert{}^{\mathrm{sc}}\Lambda_{-1/2,1/2}Q_{r}P_{V}u\rVert^{2}+\lVert E_{r}u\rVert^{2}+\lvert\langle F^{\prime\prime}_{r}u,u\rangle\rvert\,.$
The right-hand side is bounded as $r\to 0$ and since $B_{r}\to B$ as $r\to 0$, we conclude that $Bu\in L^{2}$ and
$\lVert Bu\rVert^{2}\lesssim\lVert{}^{\mathrm{sc}}\Lambda_{-1/2,1/2}QP_{V}u\rVert^{2}+\lVert Eu\rVert^{2}+\lvert\langle F^{\prime\prime}u,u\rangle\rvert\,.$
Set
$B_{1}$ $=\operatorname{Op}_{L}(x^{-\ell+1/2}\rho^{-s+1/2}b)G_{\phi}\,.$
By elliptic regularity, we have that
$\langle F^{\prime\prime}u,u\rangle$ $\lesssim\lVert B_{1}u\rVert^{2}+\lVert u\rVert_{-N,-M}^{2}\,,$
$\lVert{}^{\mathrm{sc}}\Lambda_{-1/2,1/2}QP_{V}u\rVert$ $\lesssim\lVert GP_{V}u\rVert+\lVert u\rVert_{-N,-M}\,.$
Hence, we have proved the estimate
$\lVert Bu\rVert\lesssim_{N,M}\lVert B_{1}u\rVert+\lVert Eu\rVert+\lVert GP_{V}u\rVert+\lVert u\rVert_{-N,-M}\,.$
To remove the $B_{1}$ term we inductively apply the previous estimate to the error term starting with the estimate
$\lVert B_{1}u\rVert\lesssim_{N,M}\lVert B_{2}u\rVert+\lVert Eu\rVert+\lVert GP_{V}u\rVert+\lVert u\rVert_{-N,-M}\,,$
where
$B_{2}$ $=\operatorname{Op}_{L}(x^{-\ell+1}\rho^{-s+1}b)G_{\phi}\,.$
After a finite amount of steps we have that $s-k/2<-N$ and $\ell-k/2<-M$ and therefore we have that $\lVert B_{k}u\rVert\lesssim\lVert u\rVert_{-N,-M}$. ∎
@@
Let $B_{0},E_{0}$ and $G$ be as in the proposition. First, we claim that there are $B,E$ as in Lemma 6.11 such that ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(E)$ is controlled by ${}^{3\mathrm{sc}}\!\operatorname{Ell}(E_{0})$. Indeed, this is accomplished by taking $\delta>0$ small in the definition of $q$ (and hence of $B,E$) and thus ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(E)$ is small neighborhood of the backward flow out of +$\infty$+ ∞, see (6.18) and below. Thus, by standard scattering propagation in Section 2, we have that
$\lVert Eu\rVert\lesssim_{N,M}\lVert E_{0}u\rVert+\lVert GP_{V}u\rVert+\lVert u\rVert_{-N,-M}\,$
and thus by Lemma 6.11, we control $B$ in terms of $E_{0}$
$\lVert Bu\rVert\lesssim_{N,M}\lVert E_{0}u\rVert+\lVert GP_{V}u\rVert+\lVert u\rVert_{-N,-M}\,.$
On the other hand, by the elliptic regularity of Corollary 5.13, for $\tilde{Q}\in{}^{\mathrm{sc}}\Psi^{s,\ell}$ with $+\infty\in\operatorname{Ell}_{\operatorname{ff}}(\tilde{Q})$ with ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(\tilde{Q})$ sufficiently small, we have
$\|\tilde{Q}(\operatorname{Id}-G_{\psi})u\|\lesssim\|GP_{V}u\|+\lVert u\rVert_{-N,-M},$
where we willingly relinquish the additional differential order of regularity that comes in Corollary 5.13. In fact, we choose $\tilde{Q}$ so that $\hat{N}_{\operatorname{ff},\ell}(\tilde{Q})(+\infty)=f(0)$ with $f$ given by (6.12) and set
$B^{\prime}=\tilde{Q}(\operatorname{Id}-G_{\psi})+B\,.$
We then have:
  1. (1)
    $B^{\prime}u\in L^{2}$ and satisfies the same estimate as $Bu$,
  2. (2)
    $+\infty\in\operatorname{Ell}_{\operatorname{ff}}(B^{\prime})$.
The second claim follows directly from (4.23) and
$\hat{N}_{\operatorname{ff},\ell}(B^{\prime})(+\infty)$ $=f(0)(1-\hat{N}_{\operatorname{ff}}(G_{\psi})(+\infty))+\hat{N}_{\operatorname{ff},\ell}(B)(+\infty)$
$=f(0)(1-\hat{N}_{\operatorname{ff}}(G_{\psi})(+\infty))+f(0)\hat{N}_{\operatorname{ff}}(G_{\psi})(+\infty)$
$=f(0)>0\,.$
The first claim follows from putting the bounds for $B$ and $\tilde{Q}(\operatorname{Id}-G_{\psi})$ together. More precisely,
(6.31) $\lVert B^{\prime}u\rVert\lesssim_{N,M}\lVert E_{0}u\rVert+\lVert GP_{V}u\rVert+\lVert u\rVert_{-N,-M}\,.$
Finally, we control $B_{0}u$ in terms of $B^{\prime}u$ and $E_{0}u$ as follows. Let $Q\in{}^{\mathrm{sc}}\Psi^{0,0}$ be such that
${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(Q)$ $\subset{}^{3\mathrm{sc}}\!\operatorname{Ell}(B^{\prime})\,,$
$\operatorname{WF}_{\operatorname{ff}}^{\prime}(\operatorname{Id}-Q)$ ==$\varnothing$.= ∅ .
Then
(6.32) $\lVert QB_{0}u\rVert\lesssim_{N,M}\lVert B^{\prime}u\rVert+\lVert u\rVert_{-N,-M}\,.$
Moreover, ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(\operatorname{Id}-Q)B_{0}$ is controlled by ${}^{3\mathrm{sc}}\!\operatorname{Ell}(B^{\prime})\cup{}^{3\mathrm{sc}}\!\operatorname{Ell}(E_{0})$ through flow lines which avoid $\mathrm{NP}$, and thus by the scattering propagation estimates
(6.33) $\lVert(\operatorname{Id}-Q)B_{0}u\rVert\lesssim_{N,M}\lVert E_{0}u\rVert+\lVert B^{\prime}u\rVert+\lVert GP_{V}u\rVert+\lVert u\rVert_{-N,-M}\,.$
Consequently, we obtain
$\lVert B_{0}u\rVert$ $\leq\lVert QB_{0}u\rVert+\lVert(\operatorname{Id}-Q)B_{0}u\rVert$
$\lesssim_{N,M}\lVert E_{0}u\rVert+\lVert B^{\prime}u\rVert+\lVert GP_{V}u\rVert+\lVert u\rVert_{-N,-M}\,$
$\lesssim_{N,M}\lVert E_{0}u\rVert+\lVert GP_{V}u\rVert+\lVert u\rVert_{-N,-M}\,.$
where in the last line we used (6.31). This proves the proposition. ∎
@@ 7. Radial point estimates over $C$
As in the previous section we assume that $V\in\rho_{\operatorname{mf}}{}^{3\mathrm{sc}}\Psi^{1,0}$ and $V_{0}\in S^{-1}(\mathbb{R}^{n}_{z})$ satisfy $V-V_{0}\in{}^{3\mathrm{sc}}\Psi^{1,-1}$ and
$(V-V^{*})/(2i)\in{}^{3\mathrm{sc}}\Psi^{0,-2}\,.$
@@
Let $\ell,\ell^{\prime}\in\mathbb{R}$ with $\ell\geq\ell^{\prime}>-1/2$ and
$\tau_{0}=\pm m\in W^{\perp}.$
Let $B_{1}\in{}^{3\mathrm{sc}}\Psi^{s,\ell^{\prime}}$, $G\in{}^{3\mathrm{sc}}\Psi^{s-1,\ell+1}$ and assume that ${}^{3\mathrm{sc}}\!\operatorname{Ell}(B_{1})\subset{}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(G)$. Then, there exists $B_{0}\in{}^{3\mathrm{sc}}\Psi^{s,\ell}$ with
$\tau_{0}\in\operatorname{Ell}_{\operatorname{ff}}(B_{0})$
and ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(B_{0})\subset{}^{3\mathrm{sc}}\!\operatorname{Ell}(B_{1})$ such that for any $M,N\in\mathbb{R}$, we have
$\lVert B_{0}u\rVert\lesssim\lVert GP_{V}u\rVert+\lVert B_{1}u\rVert+\lVert u\rVert_{-N,-M}\,,$
provided the right hand side is finite.
@@
Let $\ell\in\mathbb{R}$ with $\ell<-1/2$, and let $B_{0},E\in{}^{3\mathrm{sc}}\Psi^{s,\ell}$, $G\in{}^{3\mathrm{sc}}\Psi^{s-1,\ell+1}$, such that
@@
    ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(B_{0})\cup{}^{3\mathrm{sc}}\!\operatorname{Ell}(E)\subset{}^{3\mathrm{sc}}\!\operatorname{Ell}(G)$,
@@
    $\gamma_{3\mathrm{sc}}({}^{3\mathrm{sc}}\!\operatorname{Ell}(G))\cap(\mathcal{R}^{f}_{-}\cup\mathcal{R}^{p}_{+})=\varnothing$,
@@
    for every $\alpha\in\operatorname{Char}(P_{0})$ such that
    $\pi_{X,\tau}(\alpha)\in\pi_{X,\tau}\left(\gamma_{3\mathrm{sc}}\left({}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(B_{0})\right)\cap\operatorname{Char}(P_{0})\setminus(\mathcal{R}^{f}_{+}\cup\mathcal{R}^{p}_{-})\right),$
    there exists $s_{\alpha}\in(-\varepsilon,0)$ such that
    $\exp(s_{\alpha}{}^{\mathrm{sc}}\!H_{p})(\alpha)\in{}^{3\mathrm{sc}}\!\operatorname{Ell}(E)\,,$
    and for all $s\in[s_{\alpha},0]$,
    $\exp(s_{\alpha}{}^{\mathrm{sc}}\!H_{p})(\alpha)\in\pi_{3\mathrm{sc}}^{-1}({}^{3\mathrm{sc}}\!\operatorname{Ell}(G))\,.$
For any $M,N\in\mathbb{R}$, the estimate
$\lVert B_{0}u\rVert\lesssim\lVert GP_{V}u\rVert+\lVert Eu\rVert+\lVert u\rVert_{-N,-M}$
holds.
The same statement holds if +$\mathcal{R}^{f}_{+}\cup\mathcal{R}^{p}_{-}$ is replaced by $\mathcal{R}^{f}_{-}\cup\mathcal{R}^{p}_{+}$ with forward control instead of backward control.
@@
$\lVert R_{1}(\tau)\rVert_{\mathcal{B}(L^{2},H_{\mathrm{scl}}^{1,1}}\leq Cf(\tau)f^{\flat}(\tau)\,.$
@@
(7.5) $\begin{split}\hat{N}_{\operatorname{ff},\ell+1/2-\delta}(Q_{r})(\tau)&=r^{-\delta}\chi_{1}(\tau-m)\widehat{(G_{\phi})}_{\operatorname{ff}}(\tau)\,,\\ \hat{N}_{\operatorname{ff},\ell-\delta}(B_{r})(\tau)&=2r^{-\delta}\sqrt{\tau\kappa(\ell+1/2)}\chi_{1}(\tau-m)\widehat{(G_{\phi})}_{\operatorname{ff}}(\tau)\,,\\ \hat{N}_{\operatorname{ff},\ell-\delta}(E_{r})(\tau)&\equiv 0\,.\end{split}$
@@
There exists $F_{r}\in{}^{3\mathrm{sc}}\Psi^{-\infty,2(\ell-\delta)-1}$ such that $\operatorname{WF}_{3\mathrm{sc}}^{\prime}(F_{r})\subset\operatorname{WF}_{3\mathrm{sc}}^{\prime}(B)$ and
@@
$Q$ $\coloneqq\operatorname{Op}_{L}(x^{-(\ell+1/2)}q)G_{\phi}\,,$
$B$ $\coloneqq\operatorname{Op}_{L}(x^{-\ell}b_{0})G_{\phi}\,,$
$E$ $\coloneqq\operatorname{Op}_{L}(x^{-\ell}e)G_{\phi}\,.$
With these operators, we can now state the above and below threshold radial point estimates.
@@
Let $\ell,\ell^{\prime}\in\mathbb{R}$ with $\ell\geq\ell^{\prime}>-1/2$ and let $u\in H_{\mathrm{sc}}^{s,\ell^{\prime}}$ and assume that $x^{-1/2}Bu,x^{-1/2}QP_{V}u\in L^{2}$. Then $Bu\in L^{2}$ and we have the estimate
(7.6) $\lVert Bu\rVert\lesssim\lVert x^{-1/2}Bu\rVert+\lVert x^{-1/2}QP_{V}u\rVert+\lVert u\rVert_{-N,-M}\,.$
@@
Let $\ell\in\mathbb{R}$ with $\ell<-1/2$ and let $u\in H_{\mathrm{sc}}^{-N,-M}$ and assume that $x^{-1/2}Bu,x^{-1/2}QP_{V}u,Eu\in L^{2}$. Then $Bu\in L^{2}$ and we have the estimate
$\lVert Bu\rVert\lesssim\lVert x^{-1/2}Bu\rVert+\lVert x^{-1/2}QP_{V}u\rVert+\lVert Eu\rVert+\lVert u\rVert_{-N,-M}\,.$
@@
We take $\delta=\ell-\ell^{\prime}$, so that $Q_{r}\in{}^{3\mathrm{sc}}\Psi^{-\infty,\ell^{\prime}+1/2}$. As in (6.30), we have that
@@
$\leq\mu\lVert x^{1/2}Q_{r}u\rVert^{2}+\frac{1}{4\mu}\lVert x^{-1/2}\phi P_{V}u\rVert^{2}+\lvert\langle F_{r}^{\prime}u,u\rangle\rvert$
@@
By assumption the right-hand side is bounded uniformly as $r\to 0$, and since $B_{r}\to B$, we have that $Bu\in L^{2}$. The claimed estimate follows from a standard elliptic estimate for $F_{r}^{\prime}$:
$\lvert\langle F_{r}^{\prime}u,u\rangle\rvert\lesssim\lVert x^{-1/2}B_{r}u\rVert^{2}+\lVert u\rVert_{-N,-M}\,.$

### v2

@@ 6. Propagation estimates over $C$
$\lvert\operatorname{Im}\langle Q_{r}P_{V}u,Q_{r}u\rangle\rvert\geq-\frac{1}{4\mu}\lVert Q_{r}P_{V}u\rVert_{-1/2,1/2}^{2}-\mu\lVert Q_{r}u\rVert_{1/2,-1/2}^{2}\,.$
Since $Q_{r}$ and $B_{r}$ have the same principal symbol up to a constant factor, by Lemma 5.17 and Proposition 5.8, we have that
$\lVert Q_{r}u\rVert_{1/2,-1/2}^{2}\lesssim\lVert B_{r}u\rVert^{2}+\lVert\tilde{F}_{r}G_{\psi}u\rVert^{2}+\lVert u\rVert_{-N,-M}$
for some $\tilde{F}_{r}\in{}^{3\mathrm{sc}}\Psi^{-N-1,-M-1}$ with ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(Q_{r})\subset{}^{3\mathrm{sc}}\!\operatorname{Ell}(\tilde{F}_{r})$ and ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(\tilde{F}_{r}G_{\psi})\subset{}^{3\mathrm{sc}}\!\operatorname{Ell}(B_{1})$ and $\tilde{F}_{r}$ being uniformly bounded in ${}^{3\mathrm{sc}}\Psi^{s-1,\ell-1}$. Hence, choosing $\mu>0$ sufficiently small, we can absorb $Q_{r}u$ into $B_{r}u$, and arrive at the estimate
$\lVert B_{r}u\rVert^{2}\lesssim\lVert Q_{r}P_{V}u\rVert_{-1/2,1/2}^{2}+\lVert E_{r}u\rVert^{2}+\lVert\tilde{F}_{r}G_{\psi}u\rVert^{2}+\lvert\langle F^{\prime\prime}_{r}u,u\rangle\rvert\,.$
The right-hand side is bounded as $r\to 0$ and since $B_{r}\to B_{0}$ as $r\to 0$, we conclude that $B_{0}u\in L^{2}$ and
$\lVert B_{0}u\rVert^{2}\lesssim\lVert Q_{0}P_{V}u\rVert_{-1/2,1/2}^{2}+\lVert E_{0}u\rVert^{2}+\lVert B_{1}u\rVert_{s-1/2,\ell-1/2}+\lVert u\rVert_{-N,-M}$
using that by elliptic regularity, Theorem 4.20, we have that
$\lVert\tilde{F}_{r}G_{\psi}u\rVert^{2}+\lvert\langle F^{\prime\prime}_{r}u,u\rangle\rvert$ $\lesssim\lVert B_{1}u\rVert_{s-1/2,\ell-1/2}^{2}+\lVert u\rVert_{-N,-M}^{2}\,.$
@@
By using a partition of unity and the propagation estimates in the scattering calculus, it suffices to prove the estimate for $B,E\in{}^{3\mathrm{sc}}\Psi^{0,0}$ such that there exist $\tau_{0}\in\overline{W^{\perp}}$ with $\lvert\tau_{0}\rvert>m$ and $U\subset{}^{3\mathrm{sc}}\!\operatorname{Ell}(G)$ an arbitrary small neighborhood of $\tau_{0}$ with ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(BG_{\varphi})$ being controlled by ${}^{3\mathrm{sc}}\!\operatorname{Ell}(E)$ through $U$ and $\tau_{0}\in{}^{3\mathrm{sc}}\!\operatorname{Ell}(B)$.
We take the $B_{0},E_{0}$, and $Q_{0}$ as defined above with respect to the particular $\tau_{0}$. Since we can take ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(BG_{\varphi})$ to be arbitrary small provided that $\tau_{0}\in{}^{3\mathrm{sc}}\!\operatorname{Ell}(B)$, we have that
$\lVert BG_{\varphi}u\rVert_{s,\ell}\lesssim\lVert B_{0}u\rVert+\lVert u\rVert_{-N,-M}$
by Lemma 5.17 and Proposition 5.8.
In particular this means that ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(Q_{0})\subset{}^{3\mathrm{sc}}\!\operatorname{Ell}(G)$ and therefore, again by Lemma 5.17 and Proposition 5.8, we obtain
$\lVert{}^{\mathrm{sc}}\Lambda_{-1/2,1/2}Q_{0}P_{V}u\rVert$ $\lesssim\lVert GG_{\psi_{2}}P_{V}u\rVert_{s-1,\ell+1}+\lVert u\rVert_{-N,-M}$
and similarly,
$\lVert E_{0}u\rVert$ $\lesssim\lVert EG_{\psi_{1}}u\rVert_{s,\ell}+\lVert u\rVert_{-N,-M}\,.$
We choose $B_{1}\in{}^{3\mathrm{sc}}\Psi^{0,0}$ with
  1. (i)
    ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(BG_{\varphi})\cap{}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(Q_{0})\subset{}^{3\mathrm{sc}}\!\operatorname{Ell}(B_{1})\subset U$,
  2. (ii)
    ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(B_{1}G_{\varphi})\subset U$,
  3. (iii)
    ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(B_{1})\subset{}^{3\mathrm{sc}}\!\operatorname{Ell}(G^{\prime})\cap{}^{3\mathrm{sc}}\!\operatorname{Ell}(B^{\prime})$.
Note that we can arrange the first two conditions with a construction as in Section 5.2 and the third condition is trivial near $\mathrm{NP}$.
We obtain the estimate
$\lVert BG_{\varphi}u\rVert_{s,\ell}\lesssim\lVert EG_{\psi_{1}}u\rVert_{s,\ell}+\lVert GG_{\psi_{2}}P_{V}u\rVert_{s-1,\ell+1}+\lVert B_{1}u\rVert_{s-1/2,\ell-1/2}+\lVert u\rVert_{-N,-M}\,.$
The proof is finished by applying Proposition 5.14 to $Bu$ and $B_{1}u$,
$\lVert Bu\rVert_{s,\ell}$ $\lesssim\lVert BG_{\varphi}u\rVert_{s,\ell}+\lVert G^{\prime}P_{V}u\rVert_{s-2,\ell}+\lVert B^{\prime}u\rVert_{s-1,\ell-r}+\lVert u\rVert_{-N,-M}\,,$
$\lVert B_{1}u\rVert_{s-1/2,\ell-1/2}$ $\lesssim\lVert B_{1}G_{\varphi}u\rVert_{s-1/2,\ell-1/2}+\lVert G^{\prime}P_{V}u\rVert_{s-5/2,\ell}+\lVert B^{\prime}u\rVert_{s-3/2,\ell-r-1/2}+\lVert u\rVert_{-N,-M}$
and using an standard inductive argument to absorb the term $\lVert B_{1}G_{\varphi}u\rVert_{s-1/2,\ell-1/2}$ into $\lVert u\rVert_{-N,-M}$. ∎
@@ 7. Radial point estimates over $C$
As in the previous section we assume that $V\in\rho_{\operatorname{mf}}{}^{3\mathrm{sc}}\Psi^{1,0}$ is asymptotically static of order $r\geq 1$ and the Hamiltonian of its static part $V_{\pm}\in S^{-1}(\mathbb{R}^{n}_{z})$, $H_{V_{\pm}}=\Delta+m^{2}+V_{\pm}$ have purely absolutely continuous spectrum near $[m^{2},\infty)$. Moreover, we assume that
$V-V^{*}\in{}^{3\mathrm{sc}}\Psi^{0,-2}\,.$
@@
Let $\delta>0$ sufficiently small, $\varphi,\psi_{1},\psi_{2}\in C_{c}^{\infty}(\mathbb{R})$ with $\operatorname{supp}\varphi\subset(-\delta,\delta)$ and $\psi_{j}|_{(-\delta,\delta)}\equiv 1$, $\tau_{0}\in\{\pm m\}\subset W^{\perp}$, and $B,B_{1},G,G^{\prime},B^{\prime}\in{}^{3\mathrm{sc}}\Psi^{0,0}$ such that
  1. (1)
    ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(BG_{\varphi})$ is contained in a sufficiently small neighborhood of $\tau_{0}\in\dot{C}_{3\mathrm{sc}}[X;C]$,
  2. (2)
    $\tau_{0}\in\operatorname{Ell}_{\operatorname{ff}}(B)\cap\operatorname{Ell}_{\operatorname{ff}}(B_{1})$,
  3. (3)
    ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(BG_{\varphi})\subset{}^{3\mathrm{sc}}\!\operatorname{Ell}(B_{1})\cap{}^{3\mathrm{sc}}\!\operatorname{Ell}(G)$,
  4. (4)
    ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(B)\subset{}^{3\mathrm{sc}}\!\operatorname{Ell}(B^{\prime})\cap{}^{3\mathrm{sc}}\!\operatorname{Ell}(G^{\prime})$.
For all $M,N,s,s^{\prime},\ell,\ell^{\prime}\in\mathbb{R}$ with $\ell>\ell^{\prime}>-1/2$, $s>s^{\prime}$, and $u\in H_{\mathrm{sc}}^{-N,-M}$ with $B_{1}G_{\phi}u\in H_{\mathrm{sc}}^{s^{\prime},\ell^{\prime}}$, $GG_{\psi}P_{V}u\in H_{\mathrm{sc}}^{s-1,\ell+1}$, $G^{\prime}P_{V}u\in H_{\mathrm{sc}}^{s-2,\ell}$, and $B^{\prime}u\in H_{\mathrm{sc}}^{s-1,\ell-r}$, it follows that $Bu\in H_{\mathrm{sc}}^{s,\ell}$ and
$\lVert Bu\rVert_{s,\ell}$ $\leq C\big{(}\lVert B_{1}G_{\psi_{1}}u\rVert_{s^{\prime},\ell^{\prime}}+\lVert GG_{\psi_{2}}P_{V}u\rVert_{s-1,\ell+1}+\lVert G^{\prime}P_{V}u\rVert_{s-2,\ell}+\lVert B^{\prime}u\rVert_{s-1,\ell-r}$
[…]
@@
Let $\delta>0$ sufficiently small, $\varphi,\psi_{1},\psi_{2}\in C_{c}^{\infty}(\mathbb{R})$ with $\operatorname{supp}\varphi\subset(-\delta,\delta)$ and $\psi_{j}|_{(-\delta,\delta)}\equiv 1$, and $B,E,G,G^{\prime},B^{\prime}\in{}^{3\mathrm{sc}}\Psi^{0,0}$ such that
@@
    ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(BG_{\varphi})\cup{}^{3\mathrm{sc}}\!\operatorname{Ell}(E)\subset{}^{3\mathrm{sc}}\!\operatorname{Ell}(G)$,
@@
    ${}^{3\mathrm{sc}}\!\operatorname{Ell}(G)\cap{}^{3\mathrm{sc}}\mathcal{R}_{\operatorname{src}}=\varnothing$,
@@
    ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(B)\subset{}^{3\mathrm{sc}}\!\operatorname{Ell}(B^{\prime})\cap{}^{3\mathrm{sc}}\!\operatorname{Ell}(G^{\prime})$,
  5. (5)
    ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(BG_{\varphi})\setminus{}^{3\mathrm{sc}}\mathcal{R}_{\operatorname{snk}}$ is backward controlled by ${}^{3\mathrm{sc}}\!\operatorname{Ell}(E)$ through ${}^{3\mathrm{sc}}\!\operatorname{Ell}(G)$.
For all $M,N,s,\ell\in\mathbb{R}$ with $\ell<-1/2$ and $u\in H_{\mathrm{sc}}^{-N,-M}$ with $EG_{\psi_{1}}u\in H_{\mathrm{sc}}^{s,\ell}$, $GG_{\psi_{2}}P_{V}u\in H_{\mathrm{sc}}^{s-1,\ell+1}$, $G^{\prime}P_{V}u\in H_{\mathrm{sc}}^{s-2,\ell}$, and $B^{\prime}u\in H_{\mathrm{sc}}^{s-1,\ell-r}$, it follows that $Bu\in H_{\mathrm{sc}}^{s,\ell}$ and
$\lVert Bu\rVert_{s,\ell}$ $\leq C\big{(}\lVert EG_{\psi_{1}}u\rVert_{s,\ell}+\lVert GG_{\psi_{2}}P_{V}u\rVert_{s-1,\ell+1}+\lVert G^{\prime}P_{V}u\rVert_{s-2,\ell}+\lVert B^{\prime}u\rVert_{s-1,\ell-r}$
[…]
The same statement holds if the roles of ${}^{3\mathrm{sc}}\mathcal{R}_{\operatorname{src}}$ and ${}^{3\mathrm{sc}}\mathcal{R}_{\operatorname{snk}}$ are interchanged and forward control is used instead of backward control.
@@
$\lVert R_{1}(\tau)\rVert_{\mathcal{B}(L^{2},H_{\mathrm{scl}}^{1,1})}\leq Cf(\tau)f^{\flat}(\tau)\,.$
@@
(7.5) $\begin{split}\hat{N}_{\operatorname{ff},\ell+1/2-\delta}(Q_{r})(\tau)&=r^{-\delta}\chi_{1}(\tau-m)\hat{N}_{\operatorname{ff}}(G_{\phi})(\tau)\,,\\ \hat{N}_{\operatorname{ff},\ell-\delta}(B_{r})(\tau)&=2r^{-\delta}\sqrt{\tau\kappa(\ell+1/2)}\chi_{1}(\tau-m)\hat{N}_{\operatorname{ff}}(G_{\phi})(\tau)\,,\\ \hat{N}_{\operatorname{ff},\ell-\delta}(E_{r})(\tau)&\equiv 0\,.\end{split}$
@@
There exists $F_{r}\in{}^{3\mathrm{sc}}\Psi^{-\infty,2(\ell-\delta)-1}$ such that ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(F_{r})\subset{}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(B_{r})$ and
@@
$Q_{0}$ $\coloneqq\operatorname{Op}_{L}(x^{-(\ell+1/2)}q)G_{\phi}\,,$
$B_{0}$ $\coloneqq\operatorname{Op}_{L}(x^{-\ell}b_{0})G_{\phi}\,,$
$E_{0}$ $\coloneqq\operatorname{Op}_{L}(x^{-\ell}e)G_{\phi}\,.$
@@
Let $\ell>-1/2$. Let $B_{0}$ and $Q_{0}$ defined be as above and $B_{1}\in{}^{3\mathrm{sc}}\Psi^{s-1/2,\ell-1/2}$ such that
${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(Q_{0})\subset{}^{3\mathrm{sc}}\!\operatorname{Ell}(B_{1})\,.$
If $u\in H_{\mathrm{sc}}^{-N,-M}$ with $Q_{0}P_{V}\in H_{\mathrm{sc}}^{-1/2,1/2}$, and $B_{1}u\in H_{\mathrm{sc}}^{s-1/2,\ell-1/2}$, then $B_{0}u\in L^{2}$ and
$\lVert B_{0}u\rVert\leq C\left(\lVert Q_{0}P_{V}u\rVert_{-1/2,1/2}+\lVert B_{1}u\rVert_{s-1/2,\ell-1/2}+\lVert u\rVert_{-N,-M}\right)$
@@
Let $\ell<-1/2$. Let $B_{0},E_{0}$ and $Q_{0}$ defined be as above and $B_{1}\in{}^{3\mathrm{sc}}\Psi^{s-1/2,\ell-1/2}$ such that
${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(Q_{0})\subset{}^{3\mathrm{sc}}\!\operatorname{Ell}(B_{1})\,.$
If $u\in H_{\mathrm{sc}}^{-N,-M}$ with $E_{0}u\in L^{2}$, $Q_{0}P_{V}\in H_{\mathrm{sc}}^{-1/2,1/2}$, and $B_{1}u\in H_{\mathrm{sc}}^{s-1/2,\ell-1/2}$, then $B_{0}u\in L^{2}$ and
$\lVert B_{0}u\rVert\leq C\left(\lVert E_{0}u\rVert+\lVert Q_{0}P_{V}u\rVert_{-1/2,1/2}+\lVert B_{1}u\rVert_{s-1/2,\ell-1/2}+\lVert u\rVert_{-N,-M}\right)$
@@
We take $\delta=\ell-\ell^{\prime}$, so that $Q_{r}\in{}^{3\mathrm{sc}}\Psi^{-\infty,\ell^{\prime}+1/2}$ and since we are above threshold, we have that $\kappa=1$. As in (6.31), we have that
@@
$\leq\mu\lVert x^{1/2}Q_{r}u\rVert^{2}+\frac{1}{4\mu}\lVert x^{-1/2}Q_{r}P_{V}u\rVert^{2}+\lvert\langle F_{r}^{\prime}u,u\rangle\rvert$
@@
By assumption the right-hand side is bounded uniformly as $r\to 0$, and since $B_{r}\to B_{0}$, we have that $B_{0}u\in L^{2}$. The claimed estimate follows from a standard elliptic estimate for $F_{r}^{\prime}$:
$\lvert\langle F_{r}^{\prime}u,u\rangle\rvert\lesssim\lVert B_{1}u\rVert_{s-1/2,\ell-1/2}^{2}+\lVert u\rVert_{-N,-M}^{2}\,.$